The daily adult dosage of spironolactone for hypertension can be as much as 100 mg/day, divided into two doses. A bottle of spironolactone contains 160 identical tablets, which provide 40 days of treatment. How much spironolactone is in each tablet?
Answer #1

Solution: ordered drug = 100mg / day.

Bottle contains 160 tablets which is for 40 days of treatment.

Therefore, 160 ÷ 40 = 4 tablets per day.

The daily dose is 4 tablets per day which contains 100mg of the drug divided in two dose.

So, 4 tablets contains 100mg of drug.

Therefore,one tablet will contain = 100 ÷ 4 = 25 mg

So, 25mg of spironolactone is in each tablet.